40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 8521     Entry time: Thu May 2 00:34:57 2013
Author: Koji 
Type: Update 
Category: LSC 
Subject: locking 

- Routine alignment

Locked the arm cavties. Ran ASS. As this was not enough precise alignment for PRMI locking, Yarm alignment was re-adjusted by sliders.
Xarm was also aligned in the same way.

- OPLEV alignment

Once the arms were aligned, OPLEV spots were adjusted. For this adjustment, PRM had to be aligned and OPLEV servos needed to be turned off.

- LSC offset nulling

While Jenne was measuring the dark output of the POP PD, LSC offset nulling script was executed.

- Compensation of the POP spot size fix

As Jenne reported the POP path now has a lens and the denominator for the normalization got bigger.
To compensate this change, PRMI(sb) was locked by the same configuration as yesterday (i.e. AS55Q for MICH, REFL33I for PRCL). 
After some try and error, configuration for stable locking was found. 

PRCL
Signal source: REFL33I / Normalization POP110I x 1.00 / Trigger POP110I 80up 10down
Servo: input matrix 1.00 -> PRCL Servo FM3/4/5/6 Always ON G=+8.00
Actuator: output matrix 1.00 -> PRM

MICH
Signal source: AS55Q / Normalization POP110I x 0.01 / Trigger POP110I 80up 10down
Servo: input matrix 1.00 -> MICH Servo FM4/5 Always On G=-30
Actuator output matrix -1.00 -> ITMX / +1.00 -> ITMY

This suggests that POP110I signal is 5~6 times more than before the lens was installed. 

- SQRTing option for POP110I was implemented

The PRMI optical gain is derived from (Carrier)x(1st order Sideband) or (2nd order SB)x(1st order SB).
Here the carrier and the 2nd order sidebands are nonresonant.
Therefore the optical gain is proportional to the amplitude power recycling gain of the 1st order sidebands.
On the other hand, POP 2f signals are derived from the product of the 1st and -1st order sidebands.
This means that we should take a sqrt of the POP signals to compensate the recycling gain fluctuation.

Screenshot.png

- Locking with SQRT(POP110I)

PRCL
Signal source: REFL33I / Normalization SQRT(POP110I) x 10 / Trigger POP110I 10up 3down
Servo: input matrix 1.00 -> PRCL Servo FM3/4/5/6 Always ON G=+8.00
Actuator: output matrix 1.00 -> PRM

MICH
Signal source: AS55Q / Normalization SQRT(POP110I) x 0.1 / Trigger POP110I 10up 3down
Servo: input matrix 1.00 -> MICH Servo FM4/5 Always On G=-30
Actuator output matrix -1.00 -> ITMX / +1.00 -> ITMY

The lock seems not so different from the ones without SQRTing.

The spot was still moving in yaw direction. If I chose a correct alignment, I could minimize the modulation of the internal power
by misalignment. As you can see in the following plot.

Screenshot2.png

When the alignment was deviated from the optimum, the misalignment induced RIN was much worse although this was the longest lock I ever had with the PRMIsb. (more than 8 min)

Screenshot3.png

- Locking with other signal sources

REF55I/Q trial:

Demodulation phase was adjusted to make the difference of the peak heights for MICH maximized.
After the lock is acquired, I tried to swap the signal source at the input matrix. PRCL swapping was successful but
MICH swapping was not successfull.

It is much more hard to lock the interferometer with REFL55I compared with REFL33I.

REFL165I/Q trial:

As REFL165 PD never produced any useful signal, I tried to swap it with the BBPD used in the green setup.

- Borrowed the PD, power supply from the green setup.

- Put REFL165PD aside. Placed the BBPD in the path. The DC output was 0.8V. This corresponds to the input power of ~5mW.

- Checked the signal but it was very litte (several counts even at the maximum whitening gain).

- Decided to use the power reduction pick off to introduce much more light on the PD.
  This PO mirror is 90% reflector. Therefore I had to be careful no to fry the diode.
  Currently there are OD1.3 (x1/20) power attenuator to reduce the input power down to 6.5V (40mW).

- The resulting signal is very wiered suggesting the saturation of the PD at the RF stages.

- Probably I need to make a new PD circuit which has the high pass filter to reject other low frequency components.

ELOG V3.1.3-